# Lossy Compression with Data, Perception, and Classification Constraints

### v2

@@ I Introduction
[17] RD Binary[17] Source noise in signal restoration makes the perfect reconstruction impossible, leading to the persistence of tradeoffs even in the absence of a rate constraint. The details will be discussed in Section IV.
@@ I-A Motivations
Although existing works provide certain explanations and design guidance for DL in various tasks, theoretical analysis is still missing for many important applications. For example, reconstruction distortion is concerned jointly with classification accuracies in scenarios such as autonomous driving and face recognition in surveillance systems[23, 24]. Meanwhile, generative tasks and classification tasks are often jointly considered in the context of conditional generative adversarial nets (C-GAN)[25], InfoGAN[26], and CatGAN[27] where the loss function is usually a composition of GAN-loss and mutual information regularization. In these scenarios, there is a lack of sufficient theoretical characterization of the relationships between target inference tasks and generative tasks as well as traditional reconstruction tasks.
Furthermore, multi-task machine learning is more complex compared to single-task learning. In particular, different tasks may exhibit conflicting needs, making it crucial to explore the tradeoffs among them under limited resources[28]. Meanwhile, although several heuristic designs for multi-task loss functions have been proposed[28], no guidelines on balancing different tasks have been provided from an information-theoretical perspective. The existing IB principle has been successfully applied to single-task scenarios with a focus on feature extraction. A multi-task information bottleneck theory with joint consideration of generative tasks and even traditional restoration tasks is more desirable. Such a theory has great potential to provide insights into understanding tradeoffs in diverse tasks and offer design guidelines for multi-task learning.
@@ I-B Our Contributions
  • First, we investigate two open problems: the rate-distortion-classification (RDC) function and the rate-perception-classification (RPC) function. We derive the closed-form expressions for both RDC and RPC functions in binary and scalar Gaussian scenarios. As new results in characterizing task-relevant information, RDC and RPC functions complement the IB theory and RDP tradeoff, providing insights on balancing multiple tasks under a limited compression rate.
  • In contrast to existing work showing the existence of a tradeoff between classification and perception [21] in the signal degradation and restoration problem, our theoretical analysis shows that such a tradeoff does not exist in the RPC function. To account for this contradiction, we reveal that the source noise plays a decisive role, and show that once given a certain level of distortion, the classification-perception appears in the RPC function.
@@
The rest of this paper is organized as follows. In Section II we present the problem formulation and the proposed information rate-distortion-perception-classification function. The RDC and RPC functions are investigated in Section III. In Section IV, we discuss the decisive role of source noise in the existence of tradeoffs. Then, experimental results are presented in Section V. Finally, Section VI concludes the paper.
@@ II Problem Formulation
As shown in the Fig. 1, the process of lossy compression consists of an encoder and a decoder. Assume we have a source producing an i.i.d. sequence $X_{1},X_{2},\cdots,X_{n}\sim p(x)$.
  • The encoding function $f:\mathcal{X}^{n}\mapsto\{1,2,\cdots,2^{nR}\}$ maps the source $X^{n}$ to a message $M$ with a rate of $R$ bits.
  • The decoding function $g:\{1,2,\cdots,2^{nR}\}\mapsto\mathcal{\hat{X}}^{n}$ reproduces data $\hat{X}^{n}$ to satisfy task-oriented demands of downstream applications.
@@
Remark 1 (Optimality with strong asymptotical constraints).
The operational meaning of the proposed information RDPC function can be obtained from the results in [15] via the Poisson functional representation[33] and common randomness exists between the encoder and decoder. The proof of both converse and achievability follows from demonstrating that the constraints in (1a)-(1c) are functions of the joint distribution $p_{\hat{X},X}$, thereby making the information RDPC function a specific case of the information rate function defined in [15]. Note that asymptotical achievability in [15] is very strong in the sense that all constraints should be satisfied by each single message.
Remark 2 (Achievability with weak asymptotical constraints).
In a weak definition of achievability where the constraints are satisfied averagely (e.g., $\sum_{i}^{N}\mathbb{E}[\Delta(X_{n},\hat{X}_{n})]\leq D$ for some i.i.d process $\{X_{n}\}_{1}^{N}$ with margin $p_{X}$), the RDPC function is still achievable since the strong achievability implies the weak achievability. Meanwhile, we can establish the optimality of RDC function by proving its convexity in general (See Appendix E). For the RPC function, as demonstrated in Section III-B, it reduces to the RC case for the Bernoulli and scalar Gaussian sources, which are also convex. See
Nevertheless, this paper does not primarily focus on the operational definition of the RDPC function. Instead, its main objective is to provide insights into the tradeoffs between the rate and different tasks, as well as offer guidance for balancing diverse objectives in DL algorithms through the design of loss functions. The effectiveness of the RDPC function will be demonstrated with our experimental results in Section V.
@@ III-A1 Binary source
For binary sources, we characterize the closed-form solution for $R(D,C)$ as the following theorem.
Theorem 1.
@@
$R(D,C)=\begin{cases}H(b)-H(C_{1})\qquad\text{ for }D\geq C_{1}\text{ and }C_{1}\leq b,\\ H(b)-H(D)\qquad\text{ for }D<C_{1}\text{ and }D\leq b,\\ 0\qquad\text{ for }\min\{D,C_{1}\}>b,\end{cases}$
@@
Proof.
The main idea of this proof is to first utilize the rate-distortion theory[17] and Mrs. Gerber’s Lemma [18] to find a lower bound on the optimal rate. Then we prove that this lower bound is achievable. See the details in Appendix A. ∎
@@ III-A2 Gaussian source
Theorem 2.
Consider a Gaussian source $X\sim\mathcal{N}(\mu_{x},\sigma_{x}^{2})$ and a classification variable $S\sim\mathcal{N}(\mu_{s},\sigma_{s}^{2})$ with covariance $\text{Cov}(X,S)=\theta_{1}$. The problem is infeasible if $C<\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$. Otherwise, the information rate-distortion-classification function with MSE distortion is given by
$R(D,C)=\begin{cases}\frac{1}{2}\log\frac{\sigma_{x}^{2}}{D}\\ \ \ \ \ \ \ \ \ \text{ for }D\leq\sigma_{x}^{2}(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C})),\\ -\frac{1}{2}\log(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C}))\\ \ \ \ \ \ \ \ \ \text{ for }D>\sigma_{x}^{2}(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C})),\\ 0\ \ \ \ \ \ \text{ for }C>h(S)\quad\text{and}\quad D>\sigma_{x}^{2},\end{cases}$
@@
Figure 2: Visualization for RDC function of (a) a Bernoulli source and (b) Gaussian source. The left figures depict $R(D,C)$ function as a surface. The colored solid lines on the figure indicate the different rate levels. The right figures show $R(D,C)$ function along distortion-classification planes, which is the two-dimensional projection of contour lines of the left figures.
@@
The phenomenon differs from the case of RDP [8, 11, 34]. As proved in [11], in the Gaussian case, when $\sqrt{P}\geq\sigma_{X}-\sqrt{|\sigma_{x}^{2}-D|}$, the distortion constraint becomes the only active constraint, and the rate is solely determined by $D$. However, as $P$ decreases, both $P$ and $D$ are active, leading to a joint determination of the rate by $D$ and $P$. Here $D$ is always active since $I(X,\hat{X})=0$ is achievable for any $P$ when $D$ is not an active constraint[8]. Recently, a more general diagram of the transition curves determining the activeness of $P$ and $D$ in the RDP function has been depicted in [34].
@@ III-B1 Binary source
Theorem 3.
Consider a Bernoulli source $X$ and a classification variable $S$ with the binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq\frac{1}{2}$). If $C<H(S_{1})$, the problem is infeasible. Otherwise, the information rate-distortion-classification function with total variation (TV) divergence is given by
$R(P,C)=\begin{cases}H(b)-H(C_{1})\qquad\text{ for }H(S_{1})\leq C\leq H(S),\\ 0\qquad\text{ for }C>H(S),\end{cases}$
@@
The above theorem indicates that there is no tradeoff between perception and classification in the RPC relationship, since the rate only depends on the classification constraint. The proof of Theorem 3 establishes that it is always possible to find an optimal solution that achieves a TV divergence of zero while attaining the lower bound of the rate imposed by the classification constraint.
@@ III-B2 Gaussian source
$R(P,C)=\begin{cases}-\frac{1}{2}\log(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C})),\\ \ \ \ \ \ \ \ \ \ \text{ for }\frac{1}{2}\log(1\!-\!\rho^{2})\!+\!h(S)\leq C\leq h(S);\\ 0,\ \ \ \ \ \ \text{ for }C>h(S);\end{cases}$
@@
Remark 3.
An interesting observation is that RDC, RPC, and the previous RDP[8] exhibit distinct characteristics within each trinary relationship. Specifically, for the RDP case, the distortion constraint is always active as long as $R(D,P)>0$[8], while the activeness of perception depends on its relative value with distortion. For the RDC case, under a limited rate, distortion and classification are at odds with each other in an ‘antagonistic’ way where at most one constraint could be active at the same time. Finally, when considering the perception and classification constraints, the optimal rate solely depends on the classification and is irrelevant to the perception.
@@ IV-A Comparison of Lossy Compression and Signal Restoration
where $x_{0}^{\prime}=\frac{x_{0}-a}{|a|\sqrt{1+\sigma_{N}}}$ and $x_{0}^{\prime\prime}=\frac{x_{0}+a}{|a|\sqrt{1+\sigma_{N}}}$, and $\Phi(\cdot)$ is the cumulative distribution function (CDF) of the standard normal distribution. Note that the error rate is thus a function of $a$.
@@
Figure 6: Distortion-perception tradeoff (a), distortion-classification tradeoff (b) and perception-classification tradeoff (c) in the toy example.
@@
However, when dealing with a non-zero noise level, we observe that the objectives of minimizing MSE, error rate, and KL divergence do not always align, leading to tradeoffs between each pair of constraints. Specifically, for a linear denoiser, the optimal MSE is achieved at $a=0.67$, the minimum error rate is achieved at $a=0.50$, and the lowest KL-divergence is obtained at $a=0.81$ for $\sigma_{N}=1$. The value of optimal parameter $a$ depends on the noise level $\sigma_{N}$.
@@
$\mathbb{E}[\Delta(X,\hat{X})]$ $=1+a^{2}(1+\sigma_{N}^{2})-2a$
$=(1+\sigma_{N}^{2})(a-\frac{1}{1+\sigma_{N}^{2}})^{2}+\frac{\sigma_{N}^{2}}{1+\sigma_{N}^{2}},$
which is always greater than zero as long as $\sigma_{N}\neq 0$. In general, based on the results from lossy compression with noisy sources [35, 36], when the distortion is measured by MSE, we have
@@ IV-B RPC Given a Certain Level of Distortion
In section III-B, we found that there is no tradeoff between perception and classification in the RPC relationship The optimal rate only depends on the classification constraint. However, our subsequent simulations in the scalar Gaussian case will demonstrate that imposing a constraint on the distortion level causes a tradeoff to emerge.
@@
$R(P,C|D)=\min_{\sigma_{\hat{x}}}$ $-\frac{1}{2}\log(1-\frac{(\sigma_{x}^{2}+\sigma_{\hat{x}}^{2}-D)^{2}}{4\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})$
@@
$\!\!\!\!\!\!\!\!\!\!\!-\!\frac{1}{2}\log(1\!-\!\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{4}}\cdot\frac{(\sigma_{x}^{2}\!+\!\sigma_{\hat{x}}^{2}\!-\!D)^{2}}{4\sigma_{\hat{x}}^{2}})\!\geq\!h(S)\!-\!C.$

### v3

# Task-Oriented Lossy Compression with Data, Perception, and Classification Constraints

@@ I Introduction
[18] RD Binary[18] Source noise in signal restoration makes the perfect restoration impossible, leading to persistence of tradeoffs even without a rate constraint. The details will be discussed in Section IV.
@@ I-A Motivations
Although existing works provide certain explanations and design guidance for DL in various tasks, theoretical analysis is still missing for many important applications. For example, the development of smart cities and the Internet of Things (IoT) pose new challenges in efficiently compressing and analyzing massive data to meet diverse objectives, including pixel-level reconstruction and feature extraction[24, 25]. In particular, distortion is concerned jointly with classification accuracies in scenarios such as autonomous driving and face recognition in surveillance systems[26, 27]. Meanwhile, perceptual quality and fidelity on label information are often jointly considered in the context of conditional generative adversarial nets (C-GAN)[28], InfoGAN[29], and CatGAN[30]. The loss functions are usually a composition of GAN-loss and mutual information regularization. In these scenarios, there is a lack of sufficient theoretical characterization of the relationships between target inference tasks and generative tasks as well as traditional reconstruction tasks.
Furthermore, multi-task machine learning is more complex compared to single-task learning. In particular, different tasks may exhibit conflicting needs, making it crucial to explore the tradeoffs among them under limited resources[31]. Meanwhile, although several heuristic designs for multi-task loss functions have been proposed[31], no guidelines on balancing different tasks have been provided from an information-theoretical perspective. The existing IB principle has been successfully applied to single-task scenarios with a focus on feature extraction. A multi-task information bottleneck theory with joint consideration of generative tasks and even traditional restoration tasks is more desirable. Such a theory can provide insights into understanding tradeoffs in diverse tasks, which has wide applications in next-generation task-oriented communication[32, 33]. Meanwhile, it has great potential to offer design guidelines for multi-task learning.
@@ I-B Our Contributions
  • First, we investigate two open problems: the rate-distortion-classification (RDC) function and the rate-perception-classification (RPC) function. We derive the closed-form expressions for both RDC and RPC functions in binary and scalar Gaussian cases. Notably, both RDC and RPC functions exhibit distinct characteristics compared to previous theories. They complement IB theory and RDP tradeoff, and provide new insights and frameworks on characterizing task-relevant information.
  • In contrast to existing work showing the existence of a tradeoff between classification and perception [22] in the signal degradation and restoration problem, our theoretical analysis shows that such a tradeoff does not exist in the RPC function. To account for this contradiction, we reveal that the source noise plays a decisive role, and show that once given a certain level of distortion, the classification-perception tradeoff appears in the RPC function.
@@
The rest of this paper is organized as follows. In Section II we present the problem formulation and the proposed information rate-distortion-perception-classification function. The RDC and RPC functions are investigated in Section III. In Section IV, we discuss the decisive role of source noise in the existence of tradeoffs. Then, experimental results are presented in Section V. Finally, Section VII concludes the paper.
@@ II Problem Formulation
As shown in Fig. 1, the process of lossy compression consists of an encoder and a decoder.
  • The encoding function maps the source $X$ to a message with a rate of $R$ bits.
  • The decoding function reproduces data $\hat{X}$ to satisfy task-oriented demands of downstream applications.
@@
Remark 1 (Optimality with strong asymptotical constraints)
In the above formulation, we directly use the single letter representation and propose the information RDPC function to minimize $I(X;\hat{X})$ with diverse constraints. For multiple letters with i.i.d. sequence $X_{1},X_{2},\cdots,X_{n}\sim p(x)$, the operational meaning can be obtained from the results in [16] via the Poisson functional representation[38] and common randomness exists between the encoder and decoder. The proof of both converse and achievability follows from demonstrating that the constraints in (1a)-(1c) are functions of the joint distribution $p_{\hat{X},X}$, thereby making the information RDPC function a specific case of the information rate function defined in [16]. Note that asymptotical achievability in [16] is very strong in the sense that all constraints should be satisfied by each single message.
Remark 2 (Achievability with weak asymptotical constraints)
In a weak definition of achievability where the constraints are satisfied averagely (e.g., $\sum_{i}^{N}\mathbb{E}[\Delta(X_{n},\hat{X}_{n})]\leq D$ for some i.i.d process $\{X_{n}\}_{1}^{N}$ with margin $p_{X}$), the RDPC function is still achievable since the strong achievability implies the weak achievability. Meanwhile, we can establish the optimality of RDC function by proving its convexity in general. For the RPC function, as demonstrated in Section III-B, it reduces to the RC case for the Bernoulli and scalar Gaussian sources, which are also convex.
Nevertheless, this paper does not primarily focus on the operational definition of the RDPC function. Instead, its main objective is to provide insights into the tradeoffs between the rate and different tasks, and offer guidance for balancing diverse objectives in DL algorithms through the design of loss functions. The effectiveness of the RDPC function will be demonstrated with our experimental results in Section V.
@@ III-A1 Binary source
For binary sources and classification labels, we characterize the closed-form solution for $R(D,C)$ as the following theorem.
Theorem 1
@@
$R(D,C)=\begin{cases}H(b)-H(C_{1})\qquad&\text{ for }D\geq C_{1}\text{ and }C_{1}\leq b,\\ H(b)-H(D)\qquad&\text{ for }D<C_{1}\text{ and }D\leq b,\\ 0\qquad&\text{ for }\min\{D,C_{1}\}>b,\end{cases}$
@@
Proof:
The main idea of this proof is to first utilize the rate-distortion theory[18] and Mrs. Gerber’s Lemma [19] to find a lower bound on the optimal rate. Then we prove that this lower bound is achievable. Please refer to Appendix A for more details. ∎
@@ III-A2 Gaussian source
Theorem 2
Consider a Gaussian source $X\sim\mathcal{N}(\mu_{x},\sigma_{x}^{2})$ and a classification variable $S\sim\mathcal{N}(\mu_{s},\sigma_{s}^{2})$ with covariance $\text{Cov}(X,S)=\theta_{1}$. The RDC problem is infeasible if $C<\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$. Otherwise, the information rate-distortion-classification function with MSE distortion is given by
$R(D,C)=\begin{cases}\frac{1}{2}\log\frac{\sigma_{x}^{2}}{D}&\text{ for }D\leq\sigma_{x}^{2}(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C})),\\ -\frac{1}{2}\log(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C}))&\text{ for }D>\sigma_{x}^{2}(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C})),\\ 0&\text{ for }C>h(S)\quad\text{and}\quad D>\sigma_{x}^{2},\end{cases}$
@@
Figure 2: Visualization for RDC function of (a) a Bernoulli source and (b) a Gaussian source. The left figures depict $R(D,C)$ function as a surface. The colored solid lines on the figure indicate the different rate levels. The right figures show $R(D,C)$ function along distortion-classification planes, which is the two-dimensional projection of contour lines of the left figures.
@@
The phenomenon differs from the case of RDP [9][12][39]. As proved in [12], in the Gaussian case, when $\sqrt{P}\geq\sigma_{X}-\sqrt{|\sigma_{x}^{2}-D|}$, the distortion constraint becomes the only active constraint, and the rate is solely determined by $D$. However, as $P$ decreases, both $P$ and $D$ are active, leading to a joint determination of the rate by $D$ and $P$. Here $D$ is always active since $I(X,\hat{X})=0$ is achievable for any $P$ when $D$ is not an active constraint[9].
@@ III-B1 Binary source
Theorem 3
Consider a Bernoulli source $X$ and a classification variable $S$ with the binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq 1/2$). If $C<H(S_{1})$, the problem is infeasible. Otherwise, the information rate-perception-classification function with total variation (TV) divergence is given by
$R(P,C)=\begin{cases}H(b)-H(C_{1})\qquad&\text{ for }H(S_{1})\leq C\leq H(S),\\ 0\qquad&\text{ for }C>H(S),\end{cases}$
@@
The above theorem indicates that there is no tradeoff between perception and classification for the binary case, since the rate only depends on the classification constraint. The proof of Theorem 3 establishes that it is always possible to find an optimal solution that achieves a TV divergence of zero while attaining the lower bound of the rate imposed by the classification constraint.
@@ III-B2 Gaussian source
$R(P,C)=\begin{cases}-\frac{1}{2}\log(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C})),&\text{ for }\frac{1}{2}\log(1\!-\!\rho^{2})\!+\!h(S)\leq C\leq h(S);\\ 0,&\text{ for }C>h(S);\end{cases}$
@@
Intuitively, without distortion constraint, we can generate arbitrary reconstructions with the same distribution as the source (even independent of the original data) to obtain perfect perception. Conditional entropy constraint requires preserving the covariance of $S$ and $X$, which intuitively does not contradict preserving the distribution of $X$. Although our theoretical results focus on Binary and Gaussian cases, similar phenomena will be observed in experiments of real-world datasets in Section V.
Remark 3
An interesting observation is that RDC, RPC, and the previous RDP[9] exhibit distinct characteristics within each trinary relationship for both binary and Gaussian case. Specifically, for the RDP case, the distortion constraint is always active as long as $R(D,P)>0$[9], while the activeness of perception depends on its relative value with distortion. For the RDC case, under a limited rate, distortion and classification are at odds with each other in an ‘antagonistic’ way where at most one constraint could be active at the same time. Finally, when considering the perception and classification constraints, the optimal rate solely depends on the classification and is irrelevant to the perception.
Remark 4
The results derived from Theorems 1-4 can be naturally extended to $K$ classification variables if $S_{1},\cdots,S_{K}$ are independent. In the binary case, imposing multiple constraints $H(\hat{S}_{k}|\hat{X})\leq C_{k}$ for $k=1,2,\cdots,K$ results in different bounds on $H(X|\hat{X})\leq H(\tilde{C}_{k})$, where $\tilde{C}_{k}=\frac{H^{-1}(C_{k})-p_{1}}{1-2p_{1}}$ according to Mrs. Gerber’s Lemma. It is equivalent to take the most stringent constraint $\tilde{C}\triangleq\min\{\tilde{C}_{1},\cdots,\tilde{C}_{k}\}$ and determine the lower bound of $I(X,\hat{X})$. Then, employing analogous techniques as in the proofs of Theorems 1 and 3 yields similar outcomes as those for a single label. Likewise, we can also establish the equivalence between multiple constraints and the tightest constraint in the Gaussian case, so that Theorems 2 and 3 can be extended to multiple labels. Exploring scenarios involving correlated classification variables is regarded as a future research direction.
@@ IV-A Comparison of Lossy Compression and Signal Restoration
where $x_{0}^{\prime}=\frac{x_{0}-a}{|a|\sqrt{1+\sigma_{N}}}$ and $x_{0}^{\prime\prime}=\frac{x_{0}+a}{|a|\sqrt{1+\sigma_{N}}}$, and $\Phi(\cdot)$ is the cumulative distribution function of the standard normal distribution. Note that the error rate is a function of $a$.
@@
Figure 6: Distortion-perception tradeoff (a), distortion-classification tradeoff (b), and perception-classification tradeoff (c) in the toy example.
@@
However, when dealing with a non-zero noise level, we observe that the objectives of minimizing MSE, error rate, and KL divergence do not always align, leading to tradeoffs between each pair of constraints. Specifically, for a linear denoiser, the optimal MSE is achieved at $a=0.67$, the minimum error rate is achieved at $a=0.50$, and the lowest KL-divergence is obtained at $a=0.81$ when $\sigma_{N}=1$. The value of optimal parameter $a$ depends on the noise level $\sigma_{N}$.
@@
$\mathbb{E}[\Delta(X,\hat{X})]$ $=1+a^{2}(1+\sigma_{N}^{2})-2a=(1+\sigma_{N}^{2})(a-\frac{1}{1+\sigma_{N}^{2}})^{2}+\frac{\sigma_{N}^{2}}{1+\sigma_{N}^{2}},$
which is always greater than zero as long as $\sigma_{N}\neq 0$. In general, based on the results from lossy compression with a noisy source [40, 41], when the distortion is measured by MSE, we have
@@ IV-B RPC Given a Certain Level of Distortion
In section III-B, we found that there is no tradeoff between perception and classification in the RPC relationship. The optimal rate only depends on the classification constraint. However, our subsequent simulations in the scalar Gaussian case will demonstrate that imposing a constraint on the distortion level causes a tradeoff to emerge.
@@
$R(P,C|D)=\min_{\sigma_{\hat{x}}}$ $-\frac{1}{2}\log\big{(}1-\frac{(\sigma_{x}^{2}+\sigma_{\hat{x}}^{2}-D)^{2}}{4\sigma_{x}^{2}\sigma_{\hat{x}}^{2}}\big{)}$
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!-\!\frac{1}{2}\log\big{(}1\!-\!\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{4}}\!\cdot\!\frac{(\sigma_{x}^{2}\!+\!\sigma_{\hat{x}}^{2}\!-\!D)^{2}}{4\sigma_{\hat{x}}^{2}}\big{)}\!\geq\!h(S)\!-\!C.$